Grade 11General Physics

Water is filled in a flask up to a height of 20 cm .the bottom of the flask is circular with radius 10 cm. if the atmospheric pressure is 1.01*10power5 pa,find the force exerted by the water on the bottom.take g=10m per second square and density of the water=1000kg per meter cube?

nitish
8 Years agoGrade 11

1 Answer

Eshan
8 Years ago
Pressure at the bottom of the flask=P_0+\rho gh
=(1.01\times 10^5+(1000)(10)(0.2) )Pa =(1.01\times 10^5+2000)Pa =1.03\times 10^5 Pa

Hence force acting at the base=P\times A=1.03\times 10^5\times \pi (0.1)^2N =3.23\times 10^3N